Sail Inc. is considering a new project. The equipment costs $42M, will be depreciated on a...

  1. Sail Inc. is considering a new project. The equipment costs $42M, will be depreciated on a straight-line basis over 3 years to a zero book value and can be sold for 5M at the end of 3 years. It will generate earnings before interest and taxes (EBIT) of 6M per year for 3 years. There is no net working capital expenditure. The tax rate is 25%.

Sail has a target debt ratio (D/V) of 40% debt and the rest from retained earnings.

  • They currently have 3 year debt that trades at a price of $970 per bond. The coupon rate is 5.95%, and coupons are annual. The face value is $1000. The before-tax cost on any new bonds will be the same as the yield to maturity on the current bonds. Any issue costs are negligible.
  • Analysts forecast a dividend of $5.25 for next year and the current price is $38 per share. The growth rate is 3%. Issue costs for common stock are 5%.

  1. Find the IRR of the project.
  2. Calculate the WACC.
  3. Based on the IRR, will they take the project? Explain.

Unless stated otherwise, compounding is annual and payments occur at the end of the period
Answer #1
0 1 2 3
Investment -42
Salvage 5
EBIT 6 6 6
Tax (25%) -1.5 -1.5 -1.5
Profits 4.5 4.5 4.5
Depreciation 14 14 14
Cash Flows -42 18.5 18.5 22.25
IRR 18.73%

Cash flows = Investment + Salvage x (1 - tax) + Profits + Depreciation

IRR can be calculate using the same function in excel or calculator given the cash flows.

rd 7.10%
re 17.54%
WACC 12.65%

Cost of debt can be calculated using I/Y function on a calculator

N = 3, PMT = 5.95% x 1000 = 59.5, PV = -970, FV = 1000

=> Compute I/Y = 7.10% = rd

Cost of equity, re = D1 / (P x (1 - f)) + g = 5.25 / (38 x (1 - 5%)) + 3% = 17.54%

WACC = wd x rd x (1 - tax) + we x re

= 40% x 7.10% x (1 - 25%) + 60% x 17.54%

= 12.65%

Yes, they will take the project as IRR > WACC.